# Large cluster formation in a simple model for globular colloidal aggregates

### v2

@@ Abstract
A simple model of fluid whose constituent particles interact via a short range attractive and long range repulsive potential is used to model the aggregation into large spherical-like clusters made up of hundreds of particles. The model can be thought of as a simplistic representation of membraneless organelles or a straightforward rendition of colloid flocculation into large spherical aggregates. We illustrate how temperature and particle density influence the cluster size distribution and affect inter- and intra-cluster dynamics. The system is shown to exhibit two well separated length and time scales, which can be tuned by the balance between repulsive and attractive forces. Interestingly, cluster aggregates at moderate/low temperatures approach cluster glassy phases whereas the clusters itself are liquid droplets. These states present a strong suppression of density fluctuations for a significant range of relatively large wavelengths, meeting the criterion of effective disordered hyperuniform materials.
@@ I Introduction
Spontaneous pattern formation and self-limiting association can be deemed a key phenomenon both in soft matter and biological physicsStradner et al. (2004). In the former, colloidal suspensions are characteristic examples in which a variety of patterns have been identified in systems as disparate as amphiphilicsCiach et al. (2013), or proteins, such as lysozimeLiu et al. (2010) or antibodiesYearley et al. (2014) (see Kovalchuk and coworkersKovalchuk et al. (2009) and references therein for a more comprehensive review of experimental examples). As a particularly interesting illustration of self-limiting aggregates, recently Sweatman and Lue Sweatman and Lue (2019) have argued that the formation of giant clusters due to the competition between attractive and repulsive interactions might well be behind the condensation of proteins/nucleic acids into membraneless organelles (MLOs) within the nuclei of eukaryotic cells Hirose et al. (2022). This is an alternative view to the “traditional” liquid-liquid phase separation picture that is common among the structural biology communityMurthy and Fawzi (2020).
@@
As to pattern formation in soft matter systems, there is a growing interest on this subject, in particular for the possibilities that non-templated thermodynamically controlled nano-patterningSear et al. (1999); Boles et al. (2016); Barad et al. (2021) offers in the field of nanotechnology. In this connection, simple models that exhibit short range attraction and long range repulsion (SALR), Andelman et al. (1987); Seul and Andelman (1995) have shown to be able to reproduce the experimental behavior in great detail. In the pioneering work of Andelman and coworkers it was predicted by means of a mean field theory that, for a sufficiently large amplitude of repulsive interactions, modulated cluster phases will emergeArcher et al. (2007). The central role of attraction/repulsion competition in the build-up of globular (MLO-like) bubble cluster, and ordered modulated phases or bicontinuous percolating clusters that span up to macroscopic sizes, has been known for more than three decadesSeul and Andelman (1995). The shape and size of the clusters and their spatial organization is directly determined by the balance between attractive and repulsive forces, as shown by the ground state calculations of Mossa et al.Mossa et al. (2004). Obviously, thermodynamic conditions (density and temperature/pressure) play a key role as well. Recently Liu and Xi Liu and Xi (2019) proposed a classification of SALR potentials in terms of the ratio between repulsive and attractive forces and particle size. Thus SALR potential are classified into type-I (range of the attraction is less than 20% of the particle size, and range of the repulsion is a fraction of the particle size), type-II (range of the attraction less than 20% of the particle size, range of repulsion is comparable or larger than particle size) and type-III (range of attraction beyond 20% of particle size and range of repulsion as in type II). In Ref. Liu and Xi (2019) it was shown that these three classes of interactions lead to specific features as to their phase behavior and cluster morphology.
@@
The phase behavior of SALR systems has been dealt with extensively using theoretical Archer et al. (2007); Archer (2008); Archer et al. (2008); Bomont et al. (2012); Bomont and Costa (2012); Sweatman et al. (2014) and simulation approaches Archer and Wilding (2007); Klix et al. (2010); Godfrin et al. (2014). A more comprehensive summary of relevant contributions in this connection can be found in recent review works Zhuang and Charbonneau (2016); Liu and Xi (2019); Ruiz-Franco and Zaccarelli (2021). It should be stressed that in all these instances, the long range repulsive component of the interaction is essential for the self-limitation of the aggregation process, preempting condensation/demixing transitions. The latter would drive the system towards spatially separate macroscopic phasesGodfrin et al. (2014) instead of the microstructuring induced by clustering.
It is worth mentioning that other alternative approaches can explain self-limitation in aggregation processes as well. Such is the case, for instance, when the presence of specific interactions in mixtures limits the growth of the aggregate due the saturation of bonding sites by one of the componentsPalaia and Šarić (2022), in parallel with the chain termination in radical polymerization reactions.
@@
Finally, as stated by Klix and coworkers Klix et al. (2010) SALR systems are intrinsically frustrated, and this also affects their dynamics, which is mostly controlled either by mesoscopic order or by metastable disorderKlix et al. (2010). These authors actually identify three dynamically arrested phases in their SALR model with long range electrostatic repulsions: a Wigner glass (moderate packing fraction/weak attractive interactions), a cluster glassy state (low packing fraction, strong attractive interactions), and a percolating gel phase (high packing fraction/strong attractive interactions). The possibility of the existence of a Wigner glass phase in disordered systems with long range electrostatic interactions was first postulated by Bose and Wilke using mode coupling theoryBosse and Wilke (1998). Dawson et al Dawson (2002) further analyzed in detail the possibility of dynamically arrested states in colloids stemming from either attractive or repulsive forces, whose ratio, as Klix et al. Klix et al. (2010) have shown, controls both the type of dynamics and topology of the glassy states. Experimental evidence of colloidal glassy states is extensive and, for instance, numerous examples can be found in the review chapter by Weitz Weitz (2011). These glassy states are particularly interesting from the technological standpoint, since they provide an avenue to manufacture a wide range of solid-like materials starting from colloidal solutions.
@@
In this article we will revisit a particularly simple model of SALR system that can yield all types of cluster phases, and which was explored in some detail two decades ago by Imperio and Reatto Imperio and Reatto (2004, 2006); Imperio et al. (2006); Imperio and Reatto (2007). The model, in which both attractive and repulsive interactions are simple exponentials, i.e. Kac interactionsKac (1959), was initially introduced in this context by Sear and coworkersSear et al. (1999) to account for the spontaneous patterning of quantum dots at the air-water interface. Later, Archer and coworkers Archer et al. (2007); Archer and Wilding (2007) used density functional theory and extensive Monte Carlo simulations to explore its phase behavior and map its corresponding phase diagram. Integral equations have also been applied to study this SALR potential in the homogeneous region Bomont and Costa (2012). In the comprehensive work of Archer and Wilding Archer and Wilding (2007), in addition to the vapor-liquid transition, two first order transitions were identified depending on the packing fraction. Namely, a phase change at high dilution between a vapor and a fluid of liquid-like spherical clusters (globular cluster phase), and another transition between a liquid and fluid of spherical voids (bubble cluster phase). In this paper we will focus solely on the first instance. Interestingly, the same Kac-potential model was later used by Meyra and coworkers Meyra et al. (2012, 2015) to explain the formation of vegetation patterns in environments with limited resources. This illustrates how SALR effective interactions can provide a qualitative (and even quantitative) explanation for spontaneous patterning over various orders of magnitude in the spatial (and to some extent temporal) dimension. In addition to the aforementioned studies in the bulk, Bores et al. Lomba et al. (2014); Bores et al. (2015) also performed simulations and theoretical studies on the influence of confinement in disordered media on the pattern formation for the same SALR model. As mentioned, of all the possible cluster phases we will address the formation of the low density globular cluster phase, with an interaction and thermodynamics conditions tuned to yield relatively large clusters. These could actually represent a crude model of MLOs. Using large scale simulations we have analyzed the cluster size distributions, the intercluster and intracluster microscopic structure and dynamics illustrating their dependence on the thermodynamics conditions. The presence of two well separated spatial and temporal scales will be made evident.
As a consequence of the long range of the net inter-cluster interactions and relatively large attractive interparticle interaction (a type-III SALR potential), at the lowest temperature/highest density studied, clusters end-up in the vicinity of a dynamically arrested state, which falls into the class of a cluster glassy phaseKlix et al. (2010). This is apparent from the small magnitude of both particle and the cluster diffusivities, as well as in the small frequency behavior of the frequency spectrum. Moreover, in these conditions, the second peak of the cluster-cluster pair distribution function exhibits a splitting which is a characteristic feature of dense amorphous materials (e.g metallic glasses Pan et al. (2011)). We will see that the intercluster structure factor displays a strong attenuation of density fluctuations for a significant region of small wavenumbers. The system presents characteristics of a stealthy disordered hyperuniform materialTorquato and Stillinger (2003); Batten et al. (2008); Morse et al. (2023), i.e. due to the significant suppression of density fluctuations for a certain region of large wavelengths, the material becomes to some extent “invisible” to the corresponding radiation probes. The presence of an additional region with some attenuation for much shorter wavelengths (reflecting intracluster correlations) indicates that, to a smaller degree, it might be possible to have some sort of “stealthiness” for higher frequencies. This might have important implications in the field of optics and materials science.
@@
The rest of the paper is sketched as follows. In the next section we will describe the SALR model and the simulation conditions. The single cluster structure (cluster size distributions, density profiles, average life times) and dynamics will be reviewed in Section IV. Two-particle and intercluster correlations are commented upon in Section V. The article is closed with a summary of our most relevant conclusions and future prospects.
@@
Figure 1: Left) SALR interaction potential as described by Eq. (3). Right) Average intercluster potential (solid curves) and cluster-particle potential (dashed curves), computed assuming an uniform average cluster density and spherical cluster radii.
@@ II Model and methods
Figure 2: High temperature structure factor from our HNC calculations (lines) vs MD results (symbols). In these conditions the system is in the homogeneous phase displaying the initial formation of aggregates. A vertical line marks the position of $2\pi/\lambda_{0}$,
@@
Here $\sigma$ is a measure of the inner particle core diameter and will be used as unit length. Since this is a rather soft potential, the actual size, as estimated from the particle-particle distribution function is $\sigma_{eff}\approx 0.8\sigma$. As mentioned by Sear and coworkersSear et al. (1999), while the use of the attractive exponential can be justified as a more or less crude modelization of dispersive forces as in the classical works by Kac et al.Kac (1959); Kac et al. (1963), the repulsive exponential is used solely for the sake of computational simplicity. Thus, with $K_{a},K_{r}>0$, and $\alpha_{a}>\alpha_{r}>0$ one has the characteristics of a SALR potential with the presence of a maximum for a distance $r_{0}$ larger than that corresponding to the potential minimum, whose analytical expression reads
$r_{0}=\frac{\sigma}{\alpha_{a}-\alpha_{r}}\log\frac{\alpha_{a}K_{a}}{\alpha_{r}K_{r}}.$ (2)
The effect of the inner repulsive core on the value of $r_{0}$ is negligible. $r_{0}$ is a very significant correlation distance, that in the case of globular clusters determines roughly the cluster size. As will be explained below, due to the long range character of the interaction the average intercluster distance is twice as large, i.e. $\lambda_{0}=2r_{0}$. For lamellar phases, $\lambda_{0}$ determines the modulation distance, and in the bubble phase the average distance between the bubbles. In all cases the signature of this correlation length, $\lambda_{0}$, is the presence of high intensity peak in the structure factor, S(Q), located at $Q_{m}\approx 2\pi/\lambda_{0}$. For computational convenience the potential has been truncated and shifted at $R_{c}=60\sigma$, i.e., we will be dealing with an interaction given by
@@
Setting $K_{r}=1$ and $K_{a}=2$, $\sigma=4\;$Å, $\varepsilon=0.1$ kcal/mol, $\alpha_{r}=0.1$ and $\alpha_{r}=0.25$ one gets the potential depicted in the left graph of Figure 1. With this choice of parameters $r_{0}=10.7\sigma$ which, as mentioned above, determines the cluster diameter, as will be further confirmed in the next Section when discussing the average cluster density profiles. We will consider systems at relatively low number densities in order to guarantee the formation of micellar-like phases. Two total densities will be studied, $\rho\sigma^{3}=0.025,0.05$.
The fact that the correlation length is twice the position of the potential maximum is easily understood in the case of globular clusters. In this instance one can assume a uniform effective density, $\rho_{cl}^{eff}$, inside the spherical clusters of effective radius $R_{cl}^{eff}$. These two quantities can be considered as optimization parameters in a Reverse Monte Carlo procedureMcGreevy (2001), in which the interaction potential is computed from
@@
with the objective function being the intercluster pair distribution. We have performed these calculations for the aforementioned densities. With the temperature defined as kBT/$\zeta$ in terms of the potential well depth, ($\zeta=0.068$ kcal/mol, see Figure 1), $k_{B}$ being Boltzmann’s constant, the cluster-cluster potentials have been evaluated for the lowest temperature considered, kBT/$\zeta=3$. The double volume integrals over the cluster volumes are carried out numerically using a Gauss-Legendre quadrature, with $u(r)$ given by Eqs. (1) and (3). With this procedure we end up with estimated effective cluster densities of $\rho_{cl}^{eff}\sigma^{3}=1.2$ ($\rho\sigma^{3}=0.05$) and 0.83 ($\rho\sigma^{3}=0.025$), with corresponding effective radii, $R_{cl}^{eff}/\sigma=4.5$, and 4, respectively. The results obtained from Eq. (4) are presented in the right graph of Figure 1. The dashed curves correspond to the cluster-particle interaction that is simply obtained by removing one of the effective cluster densities and one volume integration from Eq. (4). We see in the Figure that the net inter-cluster interaction is extremely repulsive, orders of magnitude more intense than the interparticle potential, and dies out approximately at $\lambda_{0}=2r_{0}$. This explains why this correlation length determines the intercluster separation, as will be explicitly reflected by the first peak of the cluster-cluster pair distribution that will be discussed in the next section. Note that this is not a consequence of the cut-off introduced in (3) for computational convenience since $R_{c}\approx 3\lambda_{0}$. It is the competition between attractive and repulsive components what cancels most of both the cluster-cluster and (to a lesser extent) the cluster-particle interactions at approximately $\lambda_{0}$. Interestingly, our spherical clusters act like large repulsive spheres of approximate diameter $\lambda_{0}$ whereas their actual diameter (the precise value depends on the definition as we will see) is $\approx\lambda_{0}/2$, i.e. the excluded volume is 23 times the average cluster volume.

### v4

# On the build-up of effective hyperuniformity from large globular colloidal aggregates

@@ Abstract
A simple three-dimensional model of a fluid whose constituent particles interact via a short range attractive and long range repulsive potential is used to model the aggregation into large spherical-like clusters made up of hundreds of particles. The model can be thought of as a straightforward rendition of colloid flocculation into large spherical aggregates. We illustrate how temperature and particle density influence the cluster size distribution and affect inter- and intra-cluster dynamics. The system is shown to exhibit two well separated length and time scales, which can be tuned by the balance between repulsive and attractive forces. Interestingly, cluster aggregates at moderate/low temperatures approach a cluster glassy phase whereas cluster particles retain a local liquid-like structure. These states present a strong suppression of density fluctuations for a significant range of relatively large wavelengths, meeting the criterion of effective disordered hyperuniform materials as far as the intercluster structure is concerned.
@@ I Introduction
Spontaneous pattern formation and self-limiting association can be deemed key phenomena both in biological and soft matter physicsStradner et al. (2004). In the former, colloidal suspensions are characteristic examples in which a variety of patterns have been identified in systems as disparate as amphiphilicsCiach et al. (2013), and proteins such as lysozimeLiu et al. (2010) or antibodiesYearley et al. (2014) (see Kovalchuk and coworkersKovalchuk et al. (2009) and references therein for a more comprehensive review of experimental examples). As a particularly interesting illustration of self-limiting aggregation, recently Sweatman and Lue Sweatman and Lue (2019) have argued that the formation of giant clusters due to the competition between attractive and repulsive interactions might well be behind the condensation of proteins/nucleic acids into membraneless organelles within the nuclei of eukaryotic cells Hirose et al. (2022). This is an alternative view to the “traditional” liquid-liquid phase separation picture that is common among the structural biology communityMurthy and Fawzi (2020).
@@
As to pattern formation in soft matter systems, there is a growing interest on this subject, in particular for the possibilities that non-templated thermodynamically controlled nano-patterningSear et al. (1999); Boles et al. (2016); Barad et al. (2021) offers in the field of nanotechnology. In this connection, simple models that exhibit short range attraction and long range repulsion (SALR), Andelman et al. (1987); Seul and Andelman (1995) have shown to be able to reproduce the experimental behavior in great detail. In the pioneering work of Andelman and coworkers it was predicted by means of a mean field theory that, for a sufficiently large amplitude of repulsive interactions, modulated cluster phases will emergeArcher et al. (2007). The central role of attraction/repulsion competition in the build-up of globular bubble cluster, and ordered modulated phases or bicontinuous percolating clusters that span up to macroscopic sizes, has been known for more than three decadesSeul and Andelman (1995). The shape and size of the clusters and their spatial organization is directly determined by the balance between attractive and repulsive forces, as shown by the ground state calculations of Mossa et al.Mossa et al. (2004). Obviously, thermodynamic conditions (density and temperature/pressure) play a key role as well. Recently Liu and Xi Liu and Xi (2019) proposed a classification of SALR potentials in terms of the ratio between repulsive and attractive forces and particle size. Thus SALR potential are classified into type-I (range of the attraction is less than 20% of the particle size, and range of the repulsion is a fraction of the particle size), type-II (range of the attraction less than 20% of the particle size, range of repulsion is comparable or larger than particle size) and type-III (range of attraction beyond 20% of particle size and range of repulsion as in type II). In Ref. Liu and Xi (2019) it was shown that these three classes of interactions lead to specific features as to their phase behavior and cluster morphology.
@@
The phase behavior of SALR systems has been dealt with extensively using mean field and density functional theory approaches Archer et al. (2007); Archer (2008); Archer et al. (2008); Bomont et al. (2012); Bomont and Costa (2012); Sweatman et al. (2014); Edelmann and Roth (2016); Pini and Parola (2017), effective field theoriesCiach (2008); Ciach et al. (2013); von Konigslow et al. (2013), and simulation approaches Archer and Wilding (2007); Klix et al. (2010); Godfrin et al. (2014); Zhuang and Charbonneau (2016); Zhuang et al. (2016). A comprehensive summary of relevant contributions in this connection can be found in recent review works Liu and Xi (2019); Ruiz-Franco and Zaccarelli (2021). It should be stressed that in all these instances, the long range repulsive component of the interaction is essential for the self-limitation of the aggregation process, preempting condensation/demixing transitions. The latter would drive the system towards spatially separate macroscopic phasesGodfrin et al. (2014) instead of the microstructuring induced by clustering. It is worth mentioning that other alternative approaches can explain self-limitation in aggregation processes as well. Such is the case, for instance, when the presence of specific interactions in mixtures limits the growth of the aggregate due the saturation of bonding sites by one of the componentsPalaia and Šarić (2022), in parallel with the chain termination in radical polymerization reactions.
@@
Finally, as stated by Klix and coworkers Klix et al. (2010) SALR systems are intrinsically frustrated, and this also affects their dynamics, which is mostly controlled either by mesoscopic order or by metastable disorderKlix et al. (2010). These authors actually identify three dynamically arrested phases in their SALR model with long range electrostatic repulsions: a Wigner glass (moderate packing fraction/weak attractive interactions), a cluster glassy state (low packing fraction, strong attractive interactions), and a percolating gel phase (high packing fraction/strong attractive interactions). The possibility of the existence of a Wigner glass phase in disordered systems with long range electrostatic interactions was first postulated by Bose and Wilke using mode coupling theoryBosse and Wilke (1998). Dawson et al Dawson (2002) further analyzed in detail the possibility of dynamically arrested states in colloids stemming from either attractive or repulsive forces, whose ratio, as Klix et al. Klix et al. (2010) have shown, controls both the type of dynamics and topology of the glassy states.
Experimental evidence of colloidal glassy states is extensive and, for instance, numerous examples can be found in the review chapter by Weitz Weitz (2011). These glassy states are particularly interesting from the technological standpoint, since they provide an avenue to manufacture a wide range of solid-like materials starting from colloidal solutions, and as will be shown in this work, they offer a feasible alternative for the fabrication of disordered hyperuniform materials. These systems are exotic states, lying in between crystals and fluidsTorquato and Stillinger (2003); Torquato (2018), i.e. while being structurally disordered, they exhibit a hidden order that, in common with crystalline phases, suppresses large scale density fluctuations. Aside from its ubiquitous occurrence in physical and biological systems (cf. see Ref. Torquato (2018) for a comprehensive review of the multiple systems in which disordered hyperuniformity is present), hyperuniform materials have been shown to display particularly interesting opticalFlorescu et al. (2009a, b); Froufe-Pérez et al. (2017); Zhou et al. (2019); Milošević et al. (2019) and acoustic propertiesRomero-García et al. (2019); Chéron et al. (2022). Manufacturing hyperuniform materials from colloidal aggregates opens multiple paths to the fine tuning of their optical/acoustical properties, since in addition to the aggregate composition (and corresponding form factor of the constituent particles), the size and/or topology of the particulate aggregates can also be controlled by chemical and physical means.
@@
In this article we will revisit a particularly simple model of colloidal SALR system that can yield all types of cluster phases, and ultimately will be shown to be capable of producing disordered effective hyperuniform states. The two dimensional version of this model was explored in detail two decades ago by Imperio and Reatto Imperio and Reatto (2004, 2006); Imperio et al. (2006); Imperio and Reatto (2007). The model, in which both attractive and repulsive interactions are simple exponentials, i.e. Kac interactionsKac (1959), was initially introduced in this context by Sear and coworkersSear et al. (1999) to account for the spontaneous patterning of quantum dots at the air-water interface. Later, Archer and coworkers Archer et al. (2007); Archer and Wilding (2007); Archer et al. (2008) used density functional theory and extensive Monte Carlo simulations in a three-dimensional version of the same model, in order to explore its phase behavior and map its corresponding phase diagram. Schwanzer and KahlSchwanzer and Kahl (2010) analyzed the competition between clustering and vapor liquid condensation, and the cluster/particle dynamicsSchwanzer et al. (2016) again in two dimensions. Interestingly, the same Kac-potential model was also used by Meyra and coworkers Meyra et al. (2012, 2015) to explain the formation of vegetation patterns in environments with limited resources. This illustrates how SALR effective interactions can provide a qualitative (and even quantitative) explanation for spontaneous patterning over various orders of magnitude in the spatial (and to some extent temporal) dimension. In addition to all these studies in the bulk, Bores et al. Lomba et al. (2014); Bores et al. (2015) performed simulations and theoretical studies on the influence of confinement into disordered media on the pattern formation for the same SALR model in two dimensions as well.
In the comprehensive work of Archer and Wilding Archer and Wilding (2007), in addition to the vapor-liquid transition, two first order transitions were identified depending on the packing fraction. Namely, a phase change at high dilution between a vapor and a fluid of liquid-like spherical clusters (globular cluster phase), and another transition between a liquid and fluid of spherical voids (bubble cluster phase). In this paper we will focus solely on the first instance, with interaction parameters and thermodynamic conditions tuned to yield relatively large clusters. Here, using large scale simulations we have analyzed the cluster size distributions, the intercluster and intracluster microscopic structure and dynamics illustrating their dependence on the thermodynamic conditions. In contrast with the two-dimensional model, in which as temperature is lowered both the intracluster and intercluster structure rapidly freeze into hexagonal arrangements, the three-dimensional system retains the liquid-like structure of the cluster droplets, and the freezing of intercluster structure is frustrated, with the system displaying features characteristics of a glass. In this work, we will exploit the long range of the net inter-cluster interactions and relatively large attractive interparticle interaction (a type-III SALR potential), to find the conditions in which clusters end-up in the vicinity of a dynamically arrested state, a cluster glassy phaseKlix et al. (2010). It will be shown that these states display a strong attenuation of density fluctuations for a significant region of small wavenumbers. The system effectively resembles a stealthy disordered hyperuniform materialTorquato and Stillinger (2003); Batten et al. (2008); Morse et al. (2023), i.e. due to the suppression of density fluctuations for a certain region of large wavelengths, the material becomes to some extent “invisible” to the corresponding radiation probes.
@@
The rest of the paper is sketched as follows. In the next section we will describe the SALR model and the simulation conditions. An analysis of the onset of clustering is presented in Section III. The single cluster structure (cluster size distributions, density profiles, average life times) and dynamics will be reviewed in Section IV. Two-particle and intercluster correlations, and the analysis of hyperuniformity are commented upon in Section V. The article is closed with a summary of our most relevant conclusions and future prospects.
@@
Figure 1: (Left) SALR interaction potential as described by Eq. (3). In the inset the Fourier transform of non-singular part of the potential, $\tilde{\phi}(Q)$, (black solid curve) is also represented. For comparison in dashed blue the corresponding transform of the untruncated potential is also plotted. In vertical lines we indicate the minima of both transforms which are the characteristic wave vectors for the truncated and untruncated interaction, namely $Q_{0}$ and $Q_{0}^{\infty}$. (Right) Average intercluster potential (solid curves) and cluster-particle potential (dashed curves), computed assuming an uniform average cluster density and spherical cluster radii.
Figure 2: Snapshot of a cluster configuration in the globular phase for $\rho\sigma^{3}=0.05$, $k_{B}T/\zeta=3$. Transparent spheres represent the excluded volume around the clusters (approx. twice the particle cluster diameter) resulting from the long range interactions. The effective cluster density inferred from the excluded volume is $\rho_{C}\lambda_{0}^{3}=0.82$, corresponding to a rather dense fluid.
@@ II Model and methods
Figure 3: High temperature structure factor from our HNC calculations (lines) vs MD results (symbols). In these conditions the system is in the homogeneous phase displaying the initial formation of aggregates. A vertical line marks the position of the characteristic wave vector, $2\pi/\lambda_{0}$, The inset depicts the medium range structure factor, wherein intracluster correlations should appear. The values close to one imply absence of correlations, i.e. ideal gas behavior.
@@
Here $\sigma$ is a measure of the inner particle core diameter and will be used as unit length. Since this is a rather soft potential, the actual size, as estimated from the particle-particle distribution function is $\sigma_{eff}\approx 0.8\sigma$. As mentioned by Sear and coworkersSear et al. (1999), while the use of the attractive exponential can be justified as a more or less crude modelization of dispersive forces as in the classical works by Kac et al.Kac (1959); Kac et al. (1963), the repulsive exponential is used solely for the sake of computational simplicity. Thus, with $K_{a},K_{r}>0$, and $\alpha_{a}>\alpha_{r}>0$ one has the characteristics of a SALR potential with the presence of a maximum for a distance $d_{m}$, larger than that corresponding to the potential minimum, and whose analytical expression reads
$d_{m}=\frac{\sigma}{\alpha_{a}-\alpha_{r}}\log\frac{\alpha_{a}K_{a}}{\alpha_{r}K_{r}}.$ (2)
The effect of the inner repulsive core on the value of $d_{m}$ is negligible. $d_{m}$ is a very significant quantity that in the case of globular clusters approximately determines the cluster size, being a rough measure of its diameter. As will be explained below, due to the long range character of the interaction, the average intercluster distance is approximately twice as large, i.e. $\lambda_{m}=2d_{m}$. For computational convenience the potential has been truncated and shifted at $R_{c}=60\sigma$, i.e., we will be dealing with an interaction given by
@@
Setting $K_{r}=1$ and $K_{a}=2$, $\sigma=4\;$Å, $\varepsilon=0.1$ kcal/mol, $\alpha_{r}=0.1$ and $\alpha_{a}=0.25$ one gets the potential depicted in the left graph of Figure 1.
Alternatively, one can resort to the characteristic wave vector of the interaction (and its corresponding correlation length) which is the absolute minimum of the Fourier transform of its non-singular part, namely, the potential without the strong short range repulsion, $\phi(r)=u(r)-\varepsilon\left(\frac{\sigma}{r}\right)^{10}$. Note that for $r>2\sigma$ $u(r)$ and $\phi(r)$ are practically identical. Its connection with the structure factor and the build up of intermediate range order can easily be understood recalling that
$\lim_{Q\rightarrow 0}S(Q)=1+\rho\tilde{h}(Q)=\left[1-\rho\tilde{c}(Q)\right]^{-1}\approx\left[\chi_{0}^{-1}+\rho\tilde{\phi}(Q)/(k_{B}T)\right]^{-1}$ (4)
where the tilde denotes a Fourier transform, $\rho$ is the particle number density, $k_{B}$ is Boltzmann’s constant, and we have made use of the Ornstein-Zernike relationHansen and McDonald (2006) that connects the pair distribution function $g(r)=h(r)-1$ with the direct correlation function, $c(r)$. Additionally, $\chi_{0}$ represents the isothermal compressibility of a reference fluid of particles interacting with the repulsive short range part of the potential ($\propto r^{-10}$ in our case). This latter quantity stems from the random phase approximation (cf Ref. Archer et al. (2008)) which gives $c(r)\approx c_{0}(r)-u(r)/k_{B}T\approx c_{0}(r)-\phi(r)/k_{B}T$, being $c_{0}(r)$ the direct correlation function of the reference fluid. For our densities of interest $\chi_{0}\approx 1$. One sees that a minimum in the potential’s Fourier transform present at $Q_{0}$ implies a maximum at $S(Q_{0})$. If the denominator in the r.h,s. of (4) vanishes, the system reaches a Lifshitz point, corresponding to the onset of infinitely long ranged modulated phases. In our case one can easily determine the position of the minimum for the untruncated interaction (1), as
$Q_{0}^{\infty}$ $=$ $\sqrt{\frac{c\alpha_{a}^{2}-\alpha_{r}^{2}}{1-c}},\;\;\;c=\left(\frac{\alpha_{r}K_{r}}{\alpha_{a}K_{a}}\right)^{1/3}$
$\lambda_{0}^{\infty}$ $=$ $2\pi/Q_{0}^{\infty},$ (5)
being $\lambda_{0}^{\infty}$ being the correlation length for the system with untruncated interaction. In the inset of the left graph of Figure 1 one sees clearly the presence of a marked minimum in the untruncated potential at $Q_{0}^{\infty}\sigma=0.253$. Due to truncation this minimum is shifted in our case to $Q_{0}\sigma=0.337$. Remarkably, the correlation value estimated from $d_{m}$ (i.e. the potential maximum), namely, $0.2932\pi/\lambda_{m}=$ 0.293 is relatively close. This is simply due to fact that the lowest order approximation for the value of both quantities (in particular for $Q_{0}^{\infty}\sigma$) is $\alpha_{a}$, (here 0.25). This only applies to the specific functional form of our interaction and our choice of parameters ($\alpha_{a}K_{a}>>\alpha_{r}K_{r}$).
Now, in lamellar phases, $\lambda_{0}$ will determine the modulation distance, and in the bubble phase the average distance between the bubbles. In all cases the signature of this correlation length, $\lambda_{0}$, is the presence of a high intensity peak in the structure factor, S(Q), located at $Q_{0}$. With our choice of parameters $d_{m}=10.7\sigma$. The corresponding estimate from the characteristic wavevector $Q_{0}$ gives a distance $d_{0}=9.32\sigma$ relatively close to the former, The cluster diameter will lie somewhere between 9 and 11 molecular diameters. This will be further confirmed in the next Section when discussing the average cluster density profiles. We will consider systems at relatively low number densities in order to guarantee the formation of micellar-like phases. Two number densities will be studied, $\rho\sigma^{3}=0.025$, and 0.05. As expected from the difference between $Q_{0}^{\infty}$ and $Q_{0}$, the truncation of the SALR potential implies that cluster sizes will be somewhat smaller than those corresponding to an untruncated potential, also the corresponding correlation lengths will also be slightly smaller.
The fact that the $\lambda_{0}$ is approximately twice the position of the potential maximum is also easily understood in terms of average cluster-cluster interactions in the case of globular clusters. In this instance one can assume a uniform effective density, $\rho_{cl}^{eff}$, inside the spherical clusters of effective radius $R_{cl}^{eff}$. These two quantities can be considered as optimization parameters in a Reverse Monte Carlo procedureMcGreevy (2001), in which the interaction potential is computed from
@@
with the objective function being the intercluster pair distribution. We have performed these calculations for the aforementioned densities. With the temperature defined as kBT/$\zeta$ in terms of the potential well depth, ($\zeta=0.068$ kcal/mol, see Figure 1), The cluster-cluster potentials have been evaluated for the lowest temperature considered, kBT/$\zeta=3$. The double volume integrals over the cluster volumes are carried out numerically using a Gauss-Legendre quadrature, with $u(r)$ given by Eqs. (1) and (3). With this procedure we end up with estimated effective cluster densities of $\rho_{cl}^{eff}\sigma^{3}=1.2$ ($\rho\sigma^{3}=0.05$) and 0.83 ($\rho\sigma^{3}=0.025$), with corresponding effective radii, $R_{cl}^{eff}/\sigma=4.5$, and 4, respectively. The results obtained from Eq. (6) are presented in the right graph of Figure 1. The dashed curves correspond to the cluster-particle interaction that is simply obtained by removing one of the effective cluster densities and one volume integration from Eq. (6). We see in the Figure that the net inter-cluster interaction is extremely repulsive, orders of magnitude more intense than the interparticle potential, and dies out at $\sim\lambda_{0}$. This explains why this quantity is closely connected with the intercluster separation, as reflected by the first peak of the cluster-cluster pair distribution that will be discussed in the next section. Our spherical clusters act like large repulsive spheres of an approximate exclusion diameter $\lambda_{0}$, whereas their actual diameter (the precise value depends on the definition as we will see) is $\approx\lambda_{0}/2$, i.e. the excluded volume is 23 times the average cluster volume. This is illustrated in Figure 2 where for a simulation snapshot the excluded volume is represented by a transparent sphere. The effective cluster density as estimated from the excluded volume is in this instance $\rho_{C}\lambda_{0}^{3}=0.82$, where $\rho_{C}$ is the number density of clusters. Note that this effective density corresponds to a rather dense fluid, in particular when compared with the very low net particle density ($\rho\sigma^{3}=0.05$).